Circular No. 6952

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

NEAR-SUN COMETS D. Biesecker, Space Applications Corporation, NASA Goddard Space

Flight Center, reports provisional data (reduced by G. V. Williams) on nine more comets found in C3 coronagraphic data. Except for the first comet (2^{o} from the sun at its closest with maximum tail length 0^{o}.5), they appear to be Kreutz sungrazers and were also observed by C2. Orbit computations are deferred until definitive astrometry can be done. Comets X/1998 G3, G4, K10 and K11 (these last two having 0^{o}.5 tails and being among the brightest Kreutz comets, with K11 following K10 along the Kreutz track 4 hr later) were discovered by K. Schenk; H2 and K9 (the latter not showing a tail) by S. Stezelberger; J2 and K8 by Biesecker; and K7 by A. Vourlidas.

1998 UT  α_{2000}  δ_{2000}
Apr. 10.526  1^{h}42.^{m}2  +12^{o}06′  X/1998 G3
10.526  1 29.9  + 5 43  X/1998 G4
29.113  2 45.5  +11 31  X/1998 H2
May 10.323  3 21.5  +14 38  X/1998 J2
16.406  3 42.0  +16 05  X/1998 K7
19.240  3 49.7  +17 07  X/1998 K8
27.742  4 23.0  +17 53  X/1998 K9
31.073  4 37.0  +17 27  X/1998 K10
30.706  4 39.1  +16 50  X/1998 K11

V4334 SAGITTARII D. K. Lynch, R. W. Russell and C. J. Rice, The Aerospace Corpora-

tion; and M. Sitko, University of Cincinnati, report infrared spectroscopy of V4334 Sgr (cf. IAUC 6322) with the Infrared Telescope Facility (+ BASS, 3′′.2-diameter aperture) on Mar. 21.56, 22.60 and May 12.5 UT: “The March spectra were identical within the formal errors (5–10 percent), indicating little or no change in 24 hr. The spectra showed a smooth, featureless continuum between 3 and 13 μm that was well fit by a grey body at 1100 K. Narrowband magnitudes L′ (3.6 μm), M′ (4.9 μm) and N′ (10.5 μm) were 3.25, 2.60 and 1.72, respectively. The spectrum is consistent with thermal emission from a shell of hot dust or a heavily reddened photosphere. Given the C-rich nature of the object (Duerbeck and Benetti 1996, Ap.J. 468, L111) and the recent dust formation (IAUC 6825), carbonaceous grains are the most likely source of emission. On May 12.50 the spectrum was still featureless but had brightened by ∼ 0.1 mag at L′, M′ and N′, and it exhibited approximately the same temperature as in March.”

1998 June 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden